OpenStudy (samara8954):

Solve for n. n = 125 n = 30 n = 6 n = 5 END STUDY SESSION

OpenStudy (samara8954):

@dtuniyants

OpenStudy (samara8954):

is the answer 6

OpenStudy (anonymous):

reduce 5/25=5/(5*5)=1/5 n/30=1/5 n*5=1*30 -------->n=6

OpenStudy (samara8954):

Solve for n. n = 270 n = 10 n = 50 n = 5

OpenStudy (anonymous):

what is the equatiom?